OpenStudy (anonymous):

You have a 3.00mL of a 500 mg/dl urea standard. How much water must be added to produce a final concentration of 200mg/dl?

OpenStudy (amistre64):

mg = ..milligrams? dl =....got no clue lol

OpenStudy (amistre64):

whats mg/dl mean?

OpenStudy (anonymous):

milligrams per deciliter

OpenStudy (amistre64):

i suspected as much... :)

OpenStudy (amistre64):

3 mL = .03 decaliters...

OpenStudy (amistre64):

500*.03 = 15 500(.03) + 0(w) = 200(.03 + w) I assume

OpenStudy (amistre64):

500 = 6 + 20w 496 = 20w 24.8 = w 24.8 decaliters of water if I did it right :)

OpenStudy (amistre64):

2480 mL of water sound right to you?

OpenStudy (anonymous):

seems a little high, since the standard was only 3.00mL

OpenStudy (amistre64):

i see an error..... I did 20 and not 200; dropped an important zero :)

OpenStudy (amistre64):

500 = 6 + 200w 496 = 200w 496/200 = w water = 2.48 dl which is 248 mL

OpenStudy (amistre64):

..still off; i see another error :)

OpenStudy (amistre64):

500*.03 = 15 15 = 6 + 200w 9 = 200w 9/200 = w 0.045 decaliters of water if I did it right this time... w = 4.5 mL ....better?

OpenStudy (anonymous):

that seems correct!

OpenStudy (amistre64):

lol ..... gonna have to recall alot of prescriptions after that :)

OpenStudy (anonymous):

i appreciate it!

OpenStudy (anonymous):

may i ask how you got the 6, in 15 = 6 + 200w?

OpenStudy (anonymous):

Let t be the amount of chemical. Then t/3.00 = 500 => t = 1500 t/x = 200 => x = 1500/200 = 7.5 mL must be added

OpenStudy (amistre64):

200*.03 = 6

OpenStudy (anonymous):

No 3.5 must be added for total of 7.5

OpenStudy (amistre64):

i figure we need a 200 solution that is a combonation of .03 and "water" 200(.03 + w) = 6 + 200w ..... not that the equation is right to begin with :)

OpenStudy (anonymous):

Shoot 4.5 added to the original 3.0 to get 7.5

OpenStudy (amistre64):

ivan..lol; was I right :)

OpenStudy (anonymous):

yep both ways work, i appreciate the help!

OpenStudy (anonymous):

This is why I suck on a test

OpenStudy (amistre64):

the math I can do....sometimes, when the moon is full and the butterflies are dancing among the daisies that is :)

OpenStudy (anonymous):

i've got a few more, i'll post questions, walking me through that made so much more sense

OpenStudy (anonymous):

You must change mine to be kosher t/(3.00+ x) = 1500 => .... x = 4.5

OpenStudy (anonymous):

Or you could introduce let new amount_new = 3.00 + x Then t/amount_new = 200 or Amount_new = 1500/200 =7.5 => x = 7.5 - 3.0 = 4.5

OpenStudy (anonymous):

Thanks!